Can someone tell me where I am going wrong with trying to solve by elimination?
8x-3y=-6
5x+6y=75
5(8x-3y=-6) 40x-15y=-30
5x+6y=75
------------
35x-9y =45
Then I am stuck! Do I divide by 35x by 35 and 45 by 35? I am so confused at the moment.

Found 2 solutions by solver91311, richwmiller:
Answer by solver91311(24713)



Your first step was correct, namely multiplying the first equation by 5, but then you need to multiply the second equation by -8. Only then can you add the two equations and eliminate the variable.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by richwmiller(17219)

It would be easier to multiply the first 8x-3y=-6 by 2 and then add the new equation to the second equation 5x+6y=75 so that you eliminate y
since -6y+6y=0